OpenStudy (anonymous):

Why is (a+b)^n+1 =(a+b)^n(a+b) =a(a+b)^n +b(a+b)^n

hartnn (hartnn):

its the property: \(x^{n+1}=x*x^n\) then distributive property is used p(q+r)=pq+pr ok?

OpenStudy (anonymous):

It's a formula.

OpenStudy (anonymous):

i don't get why (a+B)^n = a(a+b)^n +b(a+b)^n

OpenStudy (anonymous):

Just give us the question xD I mean, if there is one.

OpenStudy (anonymous):

*(a+b)^n x (a+b)

ganeshie8 (ganeshie8):

p(q+r)=pq+pr

ganeshie8 (ganeshie8):

say p = (a+b)^n

hartnn (hartnn):

(a+b)^n = a(a+b)^n +b(a+b)^n<----not true (a+b)^n+1 = a(a+b)^n +b(a+b)^n<---true

OpenStudy (anonymous):

(a+b)^n x (a+b) =a(a+b)^n +b(a+b)^n

OpenStudy (anonymous):

i don't get why that is

ganeshie8 (ganeshie8):

you want proof of distrubution property ?

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

no its okay.... ..u didnt get why distribution prperty works is it ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

basically you wanto understand why 2(3+4) = 2*3 + 2*4

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

LOL that sounds pretty cool the way you put it

ganeshie8 (ganeshie8):

idk the proof id love to see @TuringTest @KingGeorge @mukushla explain it

ganeshie8 (ganeshie8):

well let me try

ganeshie8 (ganeshie8):

2*100 = 100 + 100 right ?